# Efficient Generation of Multi-partite Entanglement between Non-local Superconducting Qubits using Classical Feedback

### v2

@@ IV Quantum Fan-Out Gate
With slight modifications to the teleportation-based CNOT, it is possible to construct another quantum gate — the unbounded quantum fan-out operation Høyer and Špalek (2005) — in constant depth with the help of MCMs and feed-forward control. The unbounded quantum fan-out (see Fig. 3a) is a powerful multi-qubit gate that enables constant depth circuits to approximate various operations with a polynomially small error Høyer and Špalek (2005), and it attracts interest in the study of quantum circuit complexity Høyer and Špalek (2005). Moreover, it can help reduce circuit depth for many different tasks if it can be implemented natively, and in general gives a linear speedup over serialized instructions Gokhale et al. (2021). For example, the Quantum Fourier Transformation (QFT) — a central component of algorithms for performing discrete logarithms and factoring of integers Shor (1994) — is typically performed in quadratic depth Coppersmith (2002), but can be approximately performed in constant depth and polynomial size using the fan-out gate Høyer and Špalek (2005). Other gates like -qubit generalized Toffoli can also be implemented exactly in constant depth with the fan-out gate Takahashi and Tani (2016), which enables the possibility for arbitrary state preparation in constant depth together with one- and two-qubit operations (at the expense of exponential ancillae overhead) Rosenthal (2023). For certain states like W-state or low-Hamming-weight Dicke states, constant-depth polynomial-size preparation can be performed with MCM even when the locality constraint on two-qubit gates is considered Buhrman et al. (2023).
@@
The quantum fan-out gate is equivalent to performing a set of serialized controlled-NOTs with one control qubit and target qubits (see Fig. 3a). However, similar to the teleportation-based CNOT introduced in the previous section, this can be implemented in constant depth by means of MCM and feed-forward control. Our circuit for constant-depth implementation of quantum fan-out gate is shown in Fig. 3b. We first prepare a three-qubit GHZ state in constant depth as the entanglement resource (we are limited to three ancillae qubits, because at four qubits we cannot maintain phase coherence; see Fig. 1). Next, we entangle the control and all of the target qubits with the GHZ state in a single layer of CNOT gates. We then measure each ancilla qubit along or . Finally, based on the measurement results, if the outcome is we apply a conditional operation to all of the target qubits, and if there are an odd number of outcomes in measurements we apply a conditional gate on the control qubit. Due to the limited number of qubits, we first use two qubits as the ancillae in the constant-depth preparation of the GHZ state and then re-initialize them (using active reset) as the data qubits for the fan-out gate. In principle, with more qubits and better connectivity, we may choose two separate sets of ancillae qubits that each serve only one purpose.
@@
To demonstrate the feasibility of the fan-out gate, we implement a controlled-NOT-NOT (CXX) between three non-local qubits in constant depth. In Fig. 3c, we benchmark the CXX gate using truth table tomography and measure a truth table fidelity of . As we have previously pointed out, our qubits suffer from measurement-induced dephasing (see Appendix E). Performing truth table tomography to some degree is immune to this effect, since the data qubits are initialized and measured in computational basis (after being used to prepare ancillae qubits in a GHZ state). However, future work on other quantum processors will explore to what extent we can successfully prepare and measure in non-computational basis states (e.g., performing a controlled-Z-Z gate), as phase coherence will be important when utilizing this gate within a larger quantum computation.
@@ V Entanglement Swapping and Teleportation
Similar to the teleportation-based CNOT protocol, long-distance state teleportation can also be implemented in constant depth. In Fig. 4, we outline a state-teleportation protocol from the perspective of quantum repeaters Jiang et al. (2009); Azuma et al. (2023), a useful design to tackle imperfections during long-distance entanglement distribution in quantum networks Mastriani (2023). In a purely unitary circuit, swapping states between two ends of a register is performed by means of a cascade of SWAP gates (or alternatively, two unbounded fan-out gates; see Fig. 4a). However, by utilizing MCMs and feed-forward control, it is possible to instead teleport the state from one end of a register to another in constant depth by a quantum repeater protocol with simultaneous entanglement swapping Jiang et al. (2009). The structure of quantum repeaters is simple: ancillae qubits are prepared in a Bell state in an alternating manner, and Bell measurements are performed on overlapping qubits from each Bell state (see Fig. 4b). The measurement of each repeater can be executed in parallel, after which conditional feed-forward operations are applied to the qubit on the opposite side of the register from the qubit whose state we wish to teleport. Throughout this process, the ancillae qubits are involved in a process known as “entanglement swapping” Zukowski et al. (1993); Pan et al. (1998); Jennewein et al. (2001), since we get a long-distance entangled state from several short-distance ones. In fact, entanglement swapping generates a Bell pair distributed on two ends of the line of ancillae qubits (see Fig. 4c), and we can then use this Bell pair for teleportation tasks and other teleportation-based operations like the CNOT we showed previously.
@@
Historically, entanglement swapping and teleportation were performed on photonic systems Schmid et al. (2009); Jin et al. (2015); Tsujimoto et al. (2018), NMR Nielsen et al. (1998), and atomic systems Barrett et al. (2004); Riebe et al. (2008). Deterministic teleportation has also been demonstrated on superconducting circuits Steffen et al. (2013), but previous demonstrations of entanglement swapping on superconducting qubits Ning et al. (2019) were not deterministic111While Ref. 64 claimed “deterministic” entanglement swapping, their protocol did not implement any feed-forward operations. Therefore, they were only able to deterministically prepare an entangled Bell state, without any active control over which state they prepared.. Here, we demonstrate both deterministic state teleportation and deterministic remote entanglement generation via MCMs and feed-forward operations. In Fig. 4d, we plot the raw outputs for teleporting , , and states from one side of our quantum processor to the other. The probability of measuring the incorrect state (calculated via the total variational distance to the ideal outcome) is 11.9%, 0.6%, and 5.6% for , , and , respectively. However, measuring the classical outputs of our teleportation protocol is not sufficient when the desire is to embed this sub-circuit within a larger unitary circuit, where maintaining coherence is important. For this purpose, in Fig. 4e we plot the Pauli transfer matrix (PTM) estimated via quantum process tomography (QPT) for the state teleportation protocol, which (in contrast to the previous protocols which had multiple data qubits and, thus, a larger chance to dephase) is feasible at the level of a single qubit. Ideally, the PTM should reflect the identity operation, where the input states are prepared on a qubit on one side of the register, and the output states are measured on a qubit on the other side of the register. While we measure a process fidelity of 0.67(1) — which is well above the classical threshold of 1/2 — the Pauli and eigenvalues are not well-preserved by the process. In the PTM formalism, the preservation of a given Pauli eigenvalue is directly affected by non-commuting errors Hashim et al. (2024). Therefore, this is further evidence that our data qubits undergo strong measurement-induced dephasing during our MCMs, even with dynamical decoupling (see Appendix E). Ultimately, this will limit the utility of the protocol within larger adaptive circuits.
@@
As previously stated, when performing state teleportation, we in fact entangle the two end ancillae qubits in a Bell state via entanglement swapping. We can therefore use this procedure as an alternative to the constant-depth GHZ state introduced in Sec. II, if entanglement is only desired between two non-local qubits. By varying the input state222The input to output mapping is: , , , ., we can deterministically generate any Bell state between the two end ancillae. To demonstrate this, we deterministically prepare and , and once again measure their coherences via parity oscillations (shown in Fig. 4f). We measure Bell state fidelities of and . While both of these are above the required threshold of 1/2 for genuine entanglement, it is clear from the contrast in the parity oscillations that a significant amount of phase coherence is lost in the process.
@@ VI Conclusion and Outlook
In this work, we have demonstrated a variety of protocols for generating non-local entangled states and entangling gates in constant depth. These protocols are based on teleportation- and measurement-based quantum computing, whereby one has access to an entangled resource state which can be measured in the middle of a circuit, and whose classical outcomes can be rapidly analyzed in real-time to determine which conditional feed-forward operations to apply to the unmeasured qubits. For superconducting qubits, which have relatively fast gate times and relatively short coherence times compared to other hardware platforms, the utility of such protocols depends on the ability to perform MCMs and implement real-time feedback well within the coherence times of the qubits. With access to MCMs and fast classical feedback, these protocols enable one to greatly reduce the circuit depth of many common tasks in quantum computing. This could be particularly beneficial in the NISQ era, because it enables longer circuit depths to be performed within the coherence time of the qubits. Moreover, such protocols have similar requirements as quantum error correction (mid-circuit measurement, real-time syndrome decoding, etc.); therefore, developing the capabilities to implement teleportation- and measurement-based protocols will be beneficial to the advancement of real-time quantum error correction.
@@
Nevertheless, the utility of such protocols for superconducting qubits is severely limited by the dephasing on idling spectator qubits during MCMs. This dephasing can result from the static coupling between neighboring qubits Mitchell et al. (2021) during the measurement of one qubit Jurcevic and Govia (2022), or it can result from any unintentional direct measurement of a qubit (e.g., due to frequency crowding of the readout resonators). Depending on the strength of the dephasing, methods such as dynamical decoupling may or may not help (see Appendix E). In future work, we plan to explore ways to protect spectator qubits during MCMs. For example, in fixed-frequency systems, the static coupling can be nulled by means of simultaneous off-resonant drives on both qubits Mitchell et al. (2021); or, in systems with tunable couplers, this term can be turned off entirely during measurement. To combat the unintentional (direct) measurement of a spectator qubit, we plan to explore the degree to which one can “cloak” the qubit from the readout cavity Lledó et al. (2023). Finally, because teleportation- and measurement-based protocols are fundamentally limited by the readout fidelity of MCMs, combining them with methods of performing readout correction of MCMs Hashim et al. (2023) would be an interesting avenue of exploration.
@@ Acknowledgements
All data are available from the corresponding author upon reasonable request.
@@ Appendix A Qubit & Readout Characterization
The readout calibration for all eight qutrits is plotted in Fig. 6, which displays the discrimination boundaries used to classify all results in this work. Qutrit state discrimination can be used to monitor leakage rates outside of the computational basis; however, this is currently limited to terminating measurements, not MCMs (see Sec. C and Fig. 6). Alternatively, qutrit state discrimination can be used for improving qubit readout fidelities via excited state promotion (ESP) Mallet et al. (2009), whereby an pulse is applied to each qubit before readout, after which all state results are reclassified as in post-processing. ESP can protect qubits against decay or readout-induced decay during readout. We utilize ESP to improve qubit readout both in MCMs and final measurements. The readout fidelities with ESP are listed in Table 4.
@@ Appendix B Gate Benchmarking
The single-qubit and two-qubit gates used in this worked are benchmarked using randomized benchmarking (RB) and cycle benchmarking (CB) Erhard et al. (2019). Single-qubit Clifford infidelities measured via RB are listed in Table 4. Two-qubit Clifford and CZ gate infidelities — measured via RB and CB, respectively — are listed in Table 4. All quoted infidelities are given in terms of the process infidelity , not the average gate infidelity . These two are equal up to a simple dimensionality factor:
@@ Appendix C Classical Control Hardware
All experiments in this work were performed using the open-source control system QubiC Xu et al. (2021, 2023a). QubiC is an FPGA-based control system for superconducting qubits developed at Lawrence Berkeley National Lab. The QubiC system used for these experiments was implemented on the Xilinx ZCU216 RFSoC (RF system-on-chip) evaluation board, and uses custom gateware for real-time pulse sequencing and synthesis.
The QubiC gateware has a bank of distributed processor cores for performing pulse sequencing, parameterization, and conditional execution (i.e., control flow) Fruitwala et al. (2024). The QubiC readout DSP (digital signal processing) chain includes on-FPGA demodulation and qubit state discrimination using a threshold mechanism. Currently, the discrimination is performed for MCMs using the in-phase () component of the integrated readout pulse. If , the discriminator returns a 0; if , the discriminator returns a 1. For this reason, all of the states are calibrated to be on the right side of , and all of the and states are calibrated to be on the left side of (see Fig. 6). These state-discriminated results can then be requested by any processor core (using a special instruction) and used as inputs to arbitrary control flow/branching decisions (e.g., a while loop or if/else code block) Fruitwala et al. (2024). The total feedback latency (not including readout time) is 150 ns. After these experiments were performed, a neural network-based readout discriminator was developed for MCMs performed on QubiC which is capable of distinguishing from Vora et al. (2024).

### v3

@@ IV Quantum Fan-Out Gate
With slight modifications to the teleportation-based CNOT, it is possible to construct another quantum gate — the unbounded quantum fan-out operation hoyer2005quantum — in constant depth with the help of MCMs and feed-forward control. The unbounded quantum fan-out [see Fig. 3(a)] is a powerful multi-qubit gate that enables constant depth circuits to approximate various operations with a polynomially small error hoyer2005quantum, and it attracts interest in the study of quantum circuit complexity hoyer2005quantum. Moreover, it can help reduce circuit depth for many different tasks if it can be implemented natively, and in general gives a linear speedup over serialized instructions gokhale2021quantum. For example, the Quantum Fourier Transformation (QFT) — a central component of algorithms for performing discrete logarithms and factoring of integers shor1994algorithms — is typically performed in quadratic depth coppersmith2002approximate, but can be approximately performed in constant depth and polynomial size using the fan-out gate hoyer2005quantum. Other gates like -qubit generalized Toffoli can also be implemented exactly in constant depth with the fan-out gate takahashi_collapse_2016, which enables the possibility for arbitrary state preparation in constant depth together with one- and two-qubit operations (at the expense of exponential ancillae overhead) rosenthal_query_2023. For certain states like W-state or low-Hamming-weight Dicke states, constant-depth polynomial-size preparation can be performed with MCM even when the locality constraint on two-qubit gates is considered buhrman_state_2024-1.
@@
The quantum fan-out gate is equivalent to performing a set of serialized controlled-NOTs with one control qubit and target qubits [see Fig. 3(a)]. However, similar to the teleportation-based CNOT introduced in the previous section, this can be implemented in constant depth by means of MCM and feed-forward control. Our circuit for constant-depth implementation of quantum fan-out gate is shown in Fig. 3(b). We first prepare a three-qubit GHZ state in constant depth as the entanglement resource (we are limited to three ancillae qubits, because at four qubits we cannot maintain phase coherence; see Fig. 1). Next, we entangle the control and all of the target qubits with the GHZ state in a single layer of CNOT gates. We then measure each ancilla qubit along or . Finally, based on the measurement results, if the outcome is we apply a conditional operation to all of the target qubits, and if there are an odd number of outcomes in measurements we apply a conditional gate on the control qubit. Due to the limited number of qubits, we first use two qubits as the ancillae in the constant-depth preparation of the GHZ state and then re-initialize them (using active reset) as the data qubits for the fan-out gate. In principle, with more qubits and better connectivity, we may choose two separate sets of ancillae qubits that each serve only one purpose.
@@
To demonstrate the feasibility of the fan-out gate, we implement a controlled-NOT-NOT (CXX) between three non-local qubits in constant depth. In Fig. 3(c), we benchmark the CXX gate using truth table tomography and measure a truth table fidelity of . As we have previously pointed out, our qubits suffer from measurement-induced dephasing (see Appendix E). Performing truth table tomography to some degree is immune to this effect, since the data qubits are initialized and measured in the computational basis (after being used to prepare ancillae qubits in a GHZ state). Consequently, we observe that the fidelity is primarily limited by decoherence and gate errors [see Fig. 3(d)]. However, future work on other quantum processors will explore to what extent we can successfully prepare and measure in non-computational basis states (e.g., performing a controlled-Z-Z gate), as phase coherence will be important when utilizing this gate within a larger quantum computation.
After the first appearance of the current work, other works also reported implementing the fan-out gate using adaptive circuits with superconducting qubits song2025constant; baumer2025measurement. Both works improve upon our results. Namely, Ref. song2025constant implements the protocol on four qubits, and demonstrate that they are not limited by measurement-induced dephasing using complete state tomography performed on their output states when their fan-out gate is applied to both computational and superposition input states (though they note that dephasing errors and MCM errors still have a significant contribution to the error budget). They report fidelities of 0.76 – 0.803, though we note that these are not directly comparable to our results, since theirs are state fidelities and ours is a truth-table fidelity, which is akin to a process fidelity. They additionally show that using real-time feed-forward incurs a larger error than simply applying Pauli-frame updates in post-processing, and that at the scale of four qubits the unitary implementation of the gate has lower error than the adaptive circuit implementation. However, Ref. baumer2025measurement goes on to show that the adaptive circuit approach outperforms the unitary circuit for qubits in a linear chain, demonstrating that such adaptive-based protocols can be beneficial even on contemporary quantum processors.
@@ V Entanglement Swapping and Teleportation
Similar to the teleportation-based CNOT protocol, long-distance state teleportation can also be implemented in constant depth. In Fig. 4, we outline a state-teleportation protocol from the perspective of quantum repeaters jiang2009quantum; azuma2023quantum, a useful design to tackle imperfections during long-distance entanglement distribution in quantum networks mastriani2023simplified. In a purely unitary circuit, swapping states between two ends of a register is performed by means of a cascade of SWAP gates [or alternatively, two unbounded fan-out gates; see Fig. 4(a)]. However, by utilizing MCMs and feed-forward control, it is possible to instead teleport the state from one end of a register to another in constant depth by a quantum repeater protocol with simultaneous entanglement swapping jiang2009quantum. The structure of quantum repeaters is simple: ancillae qubits are prepared in a Bell state in an alternating manner, and Bell measurements are performed on overlapping qubits from each Bell state [see Fig. 4(b)]. The measurement of each repeater can be executed in parallel, after which conditional feed-forward operations are applied to the qubit on the opposite side of the register from the qubit whose state we wish to teleport. Throughout this process, the ancillae qubits are involved in a process known as “entanglement swapping” zukowski1993event; pan1998experimental; jennewein2001experimental, since we get a long-distance entangled state from several short-distance ones. In fact, entanglement swapping generates a Bell pair distributed on two ends of the line of ancillae qubits [see Fig. 4(c)], and we can then use this Bell pair for teleportation tasks and other teleportation-based operations like the CNOT we showed previously.
@@
Historically, entanglement swapping and teleportation were performed on photonic systems schmid2009quantum; jin2015highly; tsujimoto2018high, NMR nielsen1998complete, and atomic systems barrett2004deterministic; riebe2008deterministic. Deterministic teleportation has also been demonstrated on superconducting circuits steffen2013deterministic, but previous demonstrations of entanglement swapping on superconducting qubits ning2019deterministic were not deterministic555While Ref. ning2019deterministic claimed “deterministic” entanglement swapping, their protocol did not implement any feed-forward operations. Therefore, they were only able to deterministically prepare an entangled Bell state, without any active control over which state they prepared.. Here, we demonstrate both deterministic state teleportation and deterministic remote entanglement generation via MCMs and feed-forward operations. In Fig. 4(d), we plot the raw outputs for teleporting , , and states from one side of our quantum processor to the other. The probability of measuring the incorrect state (calculated via the total variational distance to the ideal outcome) is 11.9%, 0.6%, and 5.6% for , , and , respectively. However, measuring the classical outputs of our teleportation protocol is not sufficient when the desire is to embed this sub-circuit within a larger unitary circuit, where maintaining coherence is important. For this purpose, in Fig. 4(e) we plot the Pauli transfer matrix (PTM) estimated via quantum process tomography (QPT) for the state teleportation protocol, which (in contrast to the previous protocols which had multiple data qubits and, thus, a larger chance to dephase) is feasible at the level of a single qubit. Ideally, the PTM should reflect the identity operation, where the input states are prepared on a qubit on one side of the register, and the output states are measured on a qubit on the other side of the register. While we measure a process fidelity of 0.67(1) — which is well above the classical threshold of 1/2 — the Pauli and eigenvalues are not well-preserved by the process. In the PTM formalism, the preservation of a given Pauli eigenvalue is directly affected by non-commuting errors hashim2024practical. Therefore, this is further evidence that our data qubits undergo strong measurement-induced dephasing during our MCMs, even with dynamical decoupling (see Appendix E). Ultimately, this will limit the utility of the protocol within larger adaptive circuits.
@@
As previously stated, when performing state teleportation, we in fact entangle the two end ancillae qubits in a Bell state via entanglement swapping. We can therefore use this procedure as an alternative to the constant-depth GHZ state introduced in Sec. II, if entanglement is only desired between two non-local qubits. By varying the input state666The input to output mapping is: , , , ., we can deterministically generate any Bell state between the two end ancillae. To demonstrate this, we deterministically prepare and , and once again measure their coherences via parity oscillations [shown in Fig. 4(f)]. We measure Bell state fidelities of and . While both of these are above the required threshold of 1/2 for genuine entanglement, it is clear from the contrast in the parity oscillations that a significant amount of phase coherence is lost in the process.
@@ VI Conclusion and Outlook
In this work, we have demonstrated a variety of protocols for generating non-local entangled states and entangling gates in constant depth. These protocols are based on teleportation- and measurement-based quantum computing, whereby one has access to an entangled resource state that can be measured in the middle of a circuit, and whose classical outcomes can be rapidly analyzed in real-time to determine which conditional feed-forward operations to apply to the unmeasured qubits. For superconducting qubits, which have relatively fast gate times and relatively short coherence times compared to other hardware platforms, the utility of such protocols depends on the ability to perform MCMs and implement real-time feedback well within the coherence times of the qubits. With access to MCMs and fast classical feedback, these protocols enable one to greatly reduce the circuit depth of many common tasks in quantum computing. This could be particularly beneficial in the NISQ era, because it enables longer circuit depths to be performed within the coherence time of the qubits. While many of the applications introduced in this work could have been implemented with MCMs and post-processed updates to the Pauli frame (e.g., GHZ state preparation), rather than real-time feedback, the ultimate goal of adaptive circuits is to embed them as sub-circuits (e.g., for state preparation) within larger unitary circuits that may contain non-Clifford gates; in such cases, Pauli frame updates in post-processing would not be possible. Moreover, such protocols have similar requirements as quantum error correction (mid-circuit measurement, real-time syndrome decoding, etc.); therefore, developing the capabilities to implement teleportation- and measurement-based protocols will be beneficial to the advancement of real-time quantum error correction. This is particularly important in the fault-tolerant regimegoogle_quantum_2025; caune_demonstrating_2024, where logical non-Clifford operations are usually implemented through magic state injection. For such operations, we will need real-time error correction using the fault-tolerant measurement outcomes on the injected ancilla, from which we will apply a conditional feedforward logical Clifford (rather than a Pauli) on the state being processed, which cannot be simply postponed to the end and corrected through a Pauli frame update.
@@
In the current era, the utility of such protocols for superconducting qubits is severely limited by dephasing on idling spectator qubits during MCMs. This dephasing can result from the static coupling between neighboring qubits mitchell2021hardware during the measurement of one qubit jurcevic2022effective, or it can result from any unintentional direct measurement of a qubit (e.g., due to frequency crowding of the readout resonators). Depending on the strength of the dephasing, methods such as dynamical decoupling may or may not help (see Appendix E). As the design of superconducting quantum processors improves (e.g., better frequency allocation of qubits and resonators), and readout becomes both faster and higher fidelity, we anticipate that the fidelity of adaptive protocols will be equal to (or better than) the equivalent unitary circuits, making them a useful tool when circuit depth is prohibitive. In future work, we plan to explore ways to protect spectator qubits during MCMs. For example, in fixed-frequency systems, the static coupling can be nulled by means of simultaneous off-resonant drives on both qubits mitchell2021hardware; or, in systems with tunable couplers, this term can be turned off entirely during measurement. To combat the unintentional (direct) measurement of a spectator qubit, we plan to explore the degree to which one can “cloak” the qubit from the readout cavity lledo2023cloaking. Finally, because teleportation- and measurement-based protocols are fundamentally limited by the readout fidelity of MCMs, combining them with methods of performing readout correction of MCMs hashim2023quasi would be an interesting avenue of exploration.
@@ Acknowledgements
The data that supports the findings of this study are available within the article.
@@ Appendix A Qubit & Readout Characterization
The readout calibration for all eight qutrits is plotted in Fig. 6, which displays the discrimination boundaries used to classify all results in this work. Qutrit state discrimination can be used to monitor leakage rates outside of the computational basis; however, this is currently limited to terminating measurements, not MCMs (see Sec. C and Fig. 6). Alternatively, qutrit state discrimination can be used for improving qubit readout fidelities via excited state promotion (ESP) mallet2009single, whereby an pulse is applied to each qubit before readout, after which all state results are reclassified as in post-processing. ESP can protect qubits against decay or readout-induced decay during readout. We utilize ESP to improve qubit readout both in MCMs and final measurements. The readout fidelities with ESP are listed in Table 4.
@@ Appendix B Gate Benchmarking
The single-qubit and two-qubit gates used in this worked are benchmarked using randomized benchmarking (RB) and cycle benchmarking (CB) erhard2019characterizing. Single-qubit Clifford infidelities measured via RB are listed in Table 4. Two-qubit Clifford and CZ gate infidelities — measured via RB and CB, respectively — are listed in Table 4. All quoted infidelities are given in terms of the process infidelity , not the average gate infidelity . These two are equal up to a simple dimensionality factor:
@@ Appendix C Classical Control Hardware
All experiments in this work were performed using the open-source control system QubiC xu2021qubic; xu2023qubic. QubiC is an FPGA-based control system for superconducting qubits developed at Lawrence Berkeley National Lab. The QubiC system used for these experiments was implemented on the Xilinx ZCU216 RFSoC (RF system-on-chip) evaluation board, and uses custom gateware for real-time pulse sequencing and synthesis.
The QubiC gateware has a bank of distributed processor cores for performing pulse sequencing, parameterization, and conditional execution (i.e., control flow) fruitwala2024distributed. The QubiC readout DSP (digital signal processing) chain includes on-FPGA demodulation and qubit state discrimination using a threshold mechanism. Currently, the discrimination is performed for MCMs using the in-phase () component of the integrated readout pulse. If , the discriminator returns a 0; if , the discriminator returns a 1. For this reason, all of the states are calibrated to be on the right side of , and all of the and states are calibrated to be on the left side of (see Fig. 6). These state-discriminated results can then be requested by any processor core (using a special instruction) and used as inputs to arbitrary control flow/branching decisions (e.g., a while loop or if/else code block) fruitwala2024distributed. The total feedback latency (not including readout time) is 150 ns. After these experiments were performed, a neural network-based readout discriminator was developed for MCMs performed on QubiC which is capable of distinguishing from vora2024ml.